# A New Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

### v1

@@ 3.1 Modelling the astrophysical population of LISA source
The formation and evolution of MBBHs and their co-evolution with host galaxies pose fundamental questions in astrophysics, with implications for understanding the regulation of star formation and the growth of galaxies [65, 66]. While active galactic nuclei (AGN), powered by growing MBBHs, are believed to play a crucial role in these processes, many aspects of MBBHs growth and AGN feedback mechanisms remain poorly understood [67, 68]. One significant challenge in studying MBBHs is related to the difficulty of observing their early evolution at high redshifts. Traditional observational methods are limited to detecting only the most luminous and rapidly growing MBBHs. GWs provide a promising avenue for probing the history of MBBHs, as their propagation through the universe is essentially unobstructed [69]. The future space-based GWs detector LISA will be particularly instrumental in this area. These binaries lie within the LISA band, and LISA’s capabilities will allow for observations deep into high redshifts, shedding light on the formation and evolution of MBBHs [70]. As a consequence, our understanding of the formation and evolution of MBBHs currently heavily relies on theoretical models. Semi-analytic models of galaxy and MBBHs formation and evolution serve as powerful tools for this[71, 72, 73, 74, 75]. These models link the hierarchical formation of dark matter halos to the evolution of galaxies and their MBBHs, capturing a wide range of galaxy formation physics [76, 77, 72, 78]. In these models, the formation of MBBHs begins with the merger of two galaxies embedded in their dark matter halos. The subsequent evolution of MBBHs proceeds through various stages, from separations of 10s-100s kpc down to a few hundred pc, and eventually to separations of $\sim$ 0.001–0.01 pc, where GWs emission drives the MBBHs to coalesce. The dynamical evolution of MBH binaries is influenced by interactions with their stellar and gaseous environments, as well as with other MBBHs through three/four-body interactions [79, 80]. However, fully hydrodynamic simulations, which incorporate the complex physics involved in MBBHs evolution, often suffer from poor resolution and are constrained to relatively small volumes, resulting in limited statistical power [81, 82, 83, 84, 85, 86, 87, 88]. Additionally, these simulations are computationally very demanding. While cosmological simulations have been employed to study the evolution of unresolved MBBHs, they generally necessitate significant post-processing and involve considerable simplifications.
The formation and evolution of MBBHs and their co-evolution with host galaxies rely heavily on several key ingredients. These include the initial mass function of the MBH seeds, the delays between galaxy mergers and MBBHs mergers. These factors introduce complexities in modeling the dynamics and growth patterns of MBBHs. The literature presents a variety of seed models, delay models, reflecting the diverse theoretical approaches to these phenomena [71, 72, 73, 74, 75, 89, 90].
@@
Seed Model: There are mainly two classes of seed models in the semi-analytical modeling of MBBHs formation: the Light-Seed (LS) and Heavy-Seed (HS) models. The LS model suggests that MBH seeds are remnants of Population III stars from high-redshift, low-metallicity environments, with masses centered around 300 solar masses and a standard deviation of 0.2 dex. These seeds are located in the most massive halos formed from the collapse of 3.5-sigma peaks in the matter density field at redshifts greater than 15, allowing for mildly super-Eddington accretion rates. In contrast, the HS model posits that MBH seeds form through the collapse of protogalactic disks due to bar instabilities, typically reaching masses of approximately $10^{5}$ solar masses, and are found in similar high-redshift halos but with accretion rates capped at the Eddington limit. The choice between these models affects the formation and evolution of MBBHs, impacting the observed event rates and characteristics of MBBHs mergers. This differentiation is crucial for understanding the astrophysical processes leading to MBBHs and their subsequent mergers [91, 92, 69, 93, 88]. Recent observations by the James Webb Space Telescope (JWST)[94] provide compelling evidence favoring the HS model [95]. JWST has detected extremely luminous quasars at redshifts greater than 7, suggesting the presence of very massive black holes already in the early universe[96]. The inferred masses of these early quasars are more consistent with the higher initial masses predicted by the HS model, as the rapid formation and growth required to reach such masses challenge the LS model’s predictions. These findings indicate that massive black hole seeds might have formed through more efficient, direct collapse mechanisms as proposed by the HS model[97].
@@
All of these factors and their combinations affect the event rate and the mass model of these MBBHs, which are crucial for understanding the range of scenarios detectable by LISA (for more details, see reference [69, 72] [^1] https://people.sissa.it/~barausse/catalogs/readme.pdf ). These elements collectively influence our predictions of the frequency and characteristics of MBBHs merger events. The actual mechanisms of MBH mergers and the corresponding mass models remain largely unknown.
The total number of compact binary coalescences per unit redshift is estimated by the equation
@@
where $\frac{dV_{c}}{dz}$ represents the comoving volume element, and $R(z)$ indicates the merger rate.
@@
To capture the wide range of possible merger rate scenarios, we parameterize the merger rate $R(z)$ using a combination of two Gaussian distributions and a power law. This approach is motivated by the need to cover all possible variations in the merger rate. Different scenarios can lead to different distributions of mergers, and a flexible parametric form allows us to represent this diversity. The Gaussian distributions help capture localized features in the merger rate, while the power law accounts for the broader trends over redshift. By configuring these components with distinct parameters, we can match a variety of theoretical models and observational constraints. Similarly, the mass model is described using variable parameters that account for different theoretical combinations of the aforementioned factors. For each unique parameter set in the merger and mass models, the predicted number of observable events varies. By integrating Equation 3.1, we calculate the total number of events, allowing us to predict the observable merger events for LISA under different theoretical scenarios.
Figure 2 displays the merger rates for three distinct MBBHs models using LISA. The Light Seed model involves light MBBHs seeds derived from Light Seed stars, considering the temporal delays between MBBHs and galaxy mergers, thereby providing a more conservative estimate. The Heavy Seed+delay model assumes heavy MBBHs seeds originating from the collapse of protogalactic disks with similar delays, while Heavy Seed omits these delays, offering an optimistic scenario for LISA event rates. Each model is fitted using two Gaussian distributions and a power law function, depicted with dashed lines. This figure also illustrates the corresponding total redshifted mass $M_{z}\equiv(1+z)(m_{1}+m_{2})$ models for these scenarios in terms of the individual masses $m_{1}$ and $m_{2}$. In addition to total mass, another crucial parameter for binary systems is the mass ratio ($q\equiv m_{1}/m_{2}$, where $m_{1}<m_{2}$). We limit our analysis only for $q=1$ in this analysis for which we expect maximum matched filtering signal to noise ratio (SNR) for the sources and also more likely to detect electromagnetic counterparts. However, it’s important to note that LISA binaries can detectable for mass ratios less than one. For q within the range of $10^{-3}$ to $1$, the probability on $q$ defined as $P(q)$ in the range of q values $0.5$ to $1$ is always at least 50% for flat ($P(q)=$ constant) and for positive power law distributions with $P(q)\propto q^{\alpha},\,\alpha>0$ [102]. This percentage decreases if we assume a power law with a negative exponent in q. To select a fraction of sources with mass ratio close to one, in this analysis we consider two scenarios, namely $75\%$ (fiducial case) and $50\%$ of the total events after applying the SNR cutoff.
The models considered here are related to different theoretical approaches to MBBHs formation and evolution, significantly affecting the predicted total mass distributions and influencing the detectability of merger events by LISA. The plots effectively demonstrate how variations in initial assumptions about MBBHs seed formation and merger delays impact the mass spectrum observable by future missions. It is important to that all of these events cannot be detected due to the limitations imposed by the sensitivity of detectors. The calculation of the matched filtering SNR is essential to determine which events are detectable.
@@ 3.2 Signal to Noise Ratio (SNR) for LISA source
The matched filtering SNR, denoted as $\rho$, is crucial for detecting GW events as it quantifies the GW signal’s strength relative to the background noise. An event is deemed detectable if its SNR surpasses a predefined threshold, $\rho_{\text{TH}}$. The optimal SNR for a GW emitted by a binary system is defined by the equation [103, 104, 105]
[…]
where $\tilde{h}(f)$ represents the Fourier transform of the GWs strain, while $S_{n}(f)$ denotes the power spectral density of the detector noise. For $S_{n}(f)$, we adopt the forms from [106], and $f_{\text{min}}$ and $f_{\text{max}}$ are the boundaries for the frequency integral. Specifically, $\tilde{h}(f)$ is expressed as
$\tilde{h}(f)=h(f)Q,$ (3.3)
with $h(f)$ denoting the strain amplitude
$h(f)=\sqrt{\frac{5\eta}{24}}\frac{(GM_{\text{chirp}})^{5/6}}{d_{L}\pi^{2/3}c^{3/2}}f^{-7/6}e^{-i\psi(f)},$ (3.4)
where $M_{\text{chirp}}$ is the chirp mass, $\eta$ is the mass ratio, $d_{L}$ is the luminosity distance to the source, $c$ is the speed of light, and $\psi(f)$ is the phase of the waveform [107]. The factor $Q$ depends on the inclination angle and the detection antenna functions $F_{+}$ and $F_{\times}$
$Q=\left[F_{+}\left(\frac{1+\cos^{2}(\iota)}{2}\right)+\iota F_{\times}\cos(\iota)\right].$ (3.5)
@@
$F_{+}$ $=\frac{1}{2}(1+\cos^{2}\theta)\cos 2\phi\cos 2\psi-\cos\theta\sin 2\phi\sin 2\psi,$ (3.6)
$F_{\times}$ $=\frac{1}{2}(1+\cos^{2}\theta)\cos 2\phi\sin 2\psi+\cos\theta\sin 2\phi\cos 2\psi,$ (3.7)
@@
where $\theta$ and $\phi$ determine the source’s location in the sky, and $\psi$ relates to the orientation of the binary system with respect to the detector [108]. Consequently, the SNR formula becomes
[…]
@@
where $\Theta^{2}\equiv 4\left(F_{+}^{2}(1+\cos^{2}i)^{2}+4F_{\times}^{2}\cos^{2}i\right)$. According to [109], averaging over many binaries, the inclination angle, and sky positions, $\Theta$ follows the distribution
$P(\Theta)=\begin{cases}5\Theta(4-\Theta)^{3}/256&\text{if }0<\Theta<4,\\ 0&\text{otherwise}.\end{cases}$ (3.9)
@@
Figure 3: The figure presents the total number of injected GW events and the number of detectable GW events as a function of redshift for three different models: Light Seed, Heavy Seed, and Heavy Seed+delay. The total observation periods are indicated in parentheses next to the total number of injected events. For the detectable events, the threshold SNR is mentioned in parentheses next to the number of detectable events, assuming that EM counterparts are detected for 75% of the events.
@@
Figure 4: In this figure, we present a scatter plot illustrating the total mass of the two components of the binary system against the redshift of the selected events, with a threshold Signal-to-Noise Ratio (SNR) of 50. The plot includes data for three different models: Light Seed, Heavy Seed, and Heavy Seed+delay, represented by square, triangle, and circular markers, respectively. The observations span 4 years for the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for 75% of the events. The color of each point represents the SNR of the event, providing an additional dimension of information.
@@
In this study, we explore various merger and mass models, acknowledging the incomplete understanding of these models. By presenting different scenarios, we illustrate how the underlying assumptions about the formation and evolution of MBBHs affect the measurements of $\mathcal{F}(z)$, utilizing LISA sources. To calculate the total number of events, we integrate Equation 3.1. We determine the number of merging events for each redshift bin by selecting a bin size of $\Delta z=0.4$. As depicted in Figure 3 for the Light Seed and Heavy Seed models, the total number of events is 158 and 147, respectively, based on 4 years of observation time up to a redshift of 6. This redshift cutoff is chosen because the detection of EM counterparts is unlikely beyond this range. For the Heavy Seed+delay model, the total number of events is 49, based on 10 years of observation time, also up to a redshift of 6. For each event, we use the inverse transform method to derive the total mass of the binary system and the parameter $\Theta$ from their respective probability distributions. Subsequently, we generate an equal mass binary system and calculate the SNR using Equation 3.8. Events with an SNR greater than the threshold are selected. Since we have only considered a mass ratio(q) of 1 while generating the binary source, we retain only 75% of these events after the SNR cutoff. Additionally, we have noted the number of events if we were to select only 50% of them. For different q distributions within the range of 1e-3 to 1, the probability of getting a q value in the range of 0.5 to 1 is always $\geq$ 50% for flat and positive power law distributions. However, this percentage decreases if we assume a power law with a negative exponent in q.
For the Light Seed model, 73 out of 158 events are detected with a threshold SNR of 50. Similarly, for the Heavy Seed model, 98 out of 147 events are detected with the same threshold. For the Heavy Seed+Delay model, 31 out of 49 events are detected with a threshold SNR of 50. These numbers correspond to the 75% selection. For the 50% selection, the numbers are 51, 67, and 22, respectively. The detection probability for the Light Seed model is lower than for the other two models because the mass distribution of the Light Seed model peaks at much lower values compared to the Heavy Seed and Heavy Seed+delay models, which reduces the signal strength and consequently the SNR. In Figure 4, we show the total mass of the two components of the binary system against the redshift of the selected events, with a threshold SNR of 50.

### v2

@@ 3.1 Modelling the astrophysical population of LISA source
The formation and evolution of MBBHs and their co-evolution with host galaxies pose fundamental questions in astrophysics, with implications for understanding the regulation of star formation and the growth of galaxies [65, 66]. While active galactic nuclei (AGN), powered by growing MBBHs, are believed to play a crucial role in these processes, many aspects of MBBHs growth and AGN feedback mechanisms remain poorly understood [67, 68]. One significant challenge in studying MBBHs is related to the difficulty of observing their early evolution at high redshifts. Traditional observational methods are limited to detecting only the most luminous and rapidly growing MBBHs. GWs provide a promising avenue for probing the history of MBBHs, as their propagation through the universe is essentially unobstructed [69]. The future space-based GWs detector LISA will be particularly instrumental in this area. These binaries lie within the LISA band, and LISA’s capabilities will allow for observations deep into high redshifts, shedding light on the formation and evolution of MBBHs [70]. As a consequence, our understanding of the formation and evolution of MBBHs currently heavily relies on theoretical models. Semi-analytic models of galaxy and MBBHs formation and evolution serve as powerful tools for this[71, 72, 73, 74, 75]. These models link the hierarchical formation of dark matter halos to the evolution of galaxies and their MBBHs, capturing a wide range of galaxy formation physics [76, 77, 72, 78]. In these models, the formation of MBBHs begins with the merger of two galaxies embedded in their dark matter halos. The subsequent evolution of MBBHs proceeds through various stages, from separations of 10s-100s kpc down to a few hundred pc, and eventually to separations of $\sim$ 0.001–0.01 pc, where GWs emission drives the MBBHs to coalesce. The dynamical evolution of MBH binaries is influenced by interactions with their stellar and gaseous environments, as well as with other MBBHs through three/four-body interactions [79, 80]. However, fully hydrodynamic simulations, which incorporate the complex physics involved in MBBHs evolution, often suffer from poor resolution and are constrained to relatively small volumes, resulting in limited statistical power [81, 82, 83, 84, 85, 86, 87, 88]. Additionally, these simulations are computationally very demanding. While cosmological simulations have been employed to study the evolution of unresolved MBBHs, they generally necessitate significant post-processing and involve considerable simplifications. The formation and evolution of MBBHs and their co-evolution with host galaxies rely heavily on several key ingredients. These include the initial mass function of the MBH seeds, the delays between galaxy mergers and MBBHs mergers. These factors introduce complexities in modeling the dynamics and growth patterns of MBBHs. The literature presents a variety of seed models, delay models, reflecting the diverse theoretical approaches to these phenomena [71, 72, 73, 74, 75, 89, 90].
@@
Seed Model: There are mainly two classes of seed models in the semi-analytical modeling of MBBHs formation: the Light-Seed (LS) and Heavy-Seed (HS) models. The LS model suggests that MBH seeds are remnants of Population III stars from high-redshift, low-metallicity environments, with masses centered around 300 solar masses and a standard deviation of 0.2 dex. These seeds are located in the most massive halos formed from the collapse of 3.5-sigma peaks in the matter density field at redshifts greater than 15, allowing for mildly super-Eddington accretion rates. In contrast, the HS model posits that MBH seeds form through the collapse of protogalactic disks due to bar instabilities, typically reaching masses of approximately $\mathrm{10^{5}}$ solar masses, and are found in similar high-redshift halos but with accretion rates capped at the Eddington limit. The choice between these models affects the formation and evolution of MBBHs, impacting the observed event rates and characteristics of MBBHs mergers. This differentiation is crucial for understanding the astrophysical processes leading to MBBHs and their subsequent mergers [91, 92, 69, 93, 88]. Recent observations by the James Webb Space Telescope (JWST)[94] provide compelling evidence favoring the HS model [95]. JWST has detected extremely luminous quasars at redshifts greater than 7, suggesting the presence of very massive black holes already in the early universe[96]. The inferred masses of these early quasars are more consistent with the higher initial masses predicted by the HS model, as the rapid formation and growth required to reach such masses challenge the LS model’s predictions. These findings indicate that massive black hole seeds might have formed through more efficient, direct collapse mechanisms as proposed by the HS model[97].
@@
All of these factors and their combinations affect the event rate and the mass model of these MBBHs, which are crucial for understanding the range of scenarios detectable by LISA (for more details, see reference [69, 72]). These elements collectively influence our predictions of the frequency and characteristics of MBBHs merger events. The actual mechanisms of MBH mergers and the corresponding mass models remain largely unknown. The total number of compact binary coalescences per unit redshift is estimated by the equation
@@
where $\mathrm{\frac{dV_{c}}{dz}}$ represents the comoving volume element, and $\mathrm{R(z)}$ indicates the merger rate. To capture the wide range of possible merger rate scenarios, we parameterize the merger rate $\mathrm{R(z)}$ using a combination of two Gaussian distributions and a power law, as described by the following equation.
$\mathrm{R(z)=A}\begin{cases}\mathrm{\exp\left(-\frac{(z-\mu_{L})^{2}}{2\cdot\sigma_{L}^{2}}\right)}&\mathrm{\text{if }z<\mu_{L}}\\[10.0pt] \mathrm{\exp\left(-\alpha\cdot(z-\mu_{L})\right)}&\mathrm{\text{if }\mu_{L}\leq z\leq\mu_{R}}\\[10.0pt] \mathrm{\exp\left(-\frac{(z-\mu_{R})^{2}}{2\cdot\sigma_{R}^{2}}\right)\cdot\exp\left(-\alpha\cdot(\mu_{R}-\mu_{L})\right)}&\mathrm{\text{if }z>\mu_{R}}\end{cases}$ (3.2)
@@
where $\mathrm{A}$ is the overall normalization, $\mathrm{\mu_{L}}$ is the mean of the left Gaussian distribution, $\mathrm{\sigma_{L}}$ is the standard deviation of the left Gaussian distribution, $\mathrm{\mu_{R}}$ is the mean of the right Gaussian distribution, $\mathrm{\sigma_{R}}$ is the standard deviation of the right Gaussian distribution, and $\mathrm{\alpha}$ is the decay parameter of the exponential function. This approach is motivated by the need to cover all possible variations in the merger rate. Different scenarios can lead to different distributions of mergers, and a flexible parametric form allows us to represent this diversity. The Gaussian distributions help capture localized features in the merger rate, while the power law accounts for the broader trends over redshift. By configuring these components with distinct parameters, we can match a variety of theoretical models and observational constraints. Similarly, the mass model is described using variable parameters that account for different theoretical combinations of the aforementioned factors. For each unique parameter set in the merger and mass models, the predicted number of observable events varies. By integrating Equation 3.1, we calculate the total number of events, allowing us to predict the observable merger events for LISA under different theoretical scenarios. Figure 2 displays the merger rates for three distinct MBBH models using LISA. The Light Seed model involves light MBBH seeds derived from Light Seed stars, considering the temporal delays between MBBHs and galaxy mergers, thereby providing a more conservative estimate. The Heavy Seed+delay model assumes heavy MBBH seeds originating from the collapse of protogalactic disks with similar delays, while Heavy Seed omits these delays, offering an optimistic scenario for LISA event rates.
Each model is fitted using two Gaussian distributions and a power law function, as defined in Equation 3.2, and is depicted with dashed lines in Figure 2. The dotted points in Figure 2 represent data from semi-analytical simulations, as referenced in[69, 72] [^1] https://people.sissa.it/~barausse/catalogs/readme.pdf. The solid lines are obtained by fitting these data using Equation 3.2, and all parameter values are listed in Table 1. This figure also illustrates the corresponding total redshifted mass $\mathrm{M_{z}\equiv(1+z)(m_{1}+m_{2})}$ models for these scenarios in terms of the individual masses $\mathrm{m_{1}}$ and $\mathrm{m_{2}}$. In addition to the total mass, another crucial parameter for binary systems is the mass ratio ($\mathrm{q\equiv m_{1}/m_{2}}$, where $\mathrm{m_{1}\leq m_{2}}$). In this analysis we are interested in measuring any deviation from GR on the cosmological scale. So, we particularly focus on the high matched filtering signal to noise ratio (SNR) events in this analysis so that the uncertainty on any inferred deviation from GR is minimal. So, we consider events above SNR of fifty and only a sub-set of high SNR events are assumed to have EM counterparts. As the events with highest matched filtering SNR are those with mass ratio ($\mathrm{q=1}$) equal to one, we limit our analysis only for $\mathrm{q=1}$ in this analysis. Also, as it is currently unclear on how many LISA sources can have EM counterparts, we obtain forecast on the parameter $\mathcal{F}(z)$ for different scenarios of number of events with EM counterpart in this analysis, ranging from about $4\%$ to $75\%$ events in four years with matched filtering SNR above fifty.
[…]
Model $\mathrm{A}$ $\mathrm{\mu_{L}}$ $\mathrm{\mu_{R}}$ $\mathrm{\sigma_{L}}$ $\mathrm{\sigma_{R}}$ $\mathrm{\alpha}$
Light Seed 0.056 9.6 13.6 4.8 1.9 0.21
Heavy Seed 0.028 5.8 13.5 5 2.3 0.04
Heavy Seed + Delay 0.0045 1.5 1.8 10.5 4.5 0.02
Table 1: Parameter values for different parameters in the models used in the merger rate $\mathrm{R(z)}$ calculations. The parameters include amplitude $\mathrm{A}$, mean $\mathrm{\mu_{L}}$ and standard deviation $\mathrm{\sigma_{L}}$ for left Gaussian, mean $\mathrm{\mu_{R}}$ and standard deviation $\mathrm{\sigma_{R}}$ for right Gaussian, and decay parameter $\mathrm{\alpha}$.
Galaxy simulations to forecast the EM counterparts for massive binary mergers typically assume a well-understood evolution of galaxies and star formation rates, calibrated from large-scale surveys. However, simulating the formation and distribution of galaxies, especially at high redshifts, comes with significant uncertainties[103, 104, 105]. One of the major challenges arises from the assumptions about gas accretion, feedback processes, and the precise merger history, which can vary significantly across different simulation models. Recent observations from the JWST have further complicated our understanding of the high-redshift universe. JWST has revealed a higher abundance of massive galaxies and earlier star formation than previously predicted, challenging existing models of galaxy formation and evolution. JWST has also detected a variety of galaxies with different dust and gas content, which directly affects the probability of producing an EM counterpart. Galaxies rich in gas are more likely to produce bright EM counterparts during binary mergers. Given these uncertainties, we assume that approximately 75% and 50% of massive binary mergers may have observable EM counterparts, depending on their redshift and environment. We have also shown the results for scenarios with $25\%$ and $10\%$ bright sirens in the appendix A. This assumption reflects the current understanding of galaxy formation processes and the likelihood that such mergers would occur in gas-rich environments, which are more conducive to generating EM radiation. While these fractions are estimates, they align with predictions from theoretical models and observational data, with the emerging JWST results suggesting a broader range of environments that could host potential EM counterparts. The models considered here are related to different theoretical approaches to MBBHs formation and evolution, significantly affecting the predicted total mass distributions and influencing the detectability of merger events by LISA. The plots effectively demonstrate how variations in initial assumptions about MBBHs seed formation and merger delays impact the mass spectrum observable by future missions. It is important to that all of these events cannot be detected due to the limitations imposed by the sensitivity of detectors. The calculation of the matched filtering SNR is essential to determine which events are detectable.
@@ 3.2 Signal to Noise Ratio (SNR) for LISA source
The matched filtering SNR, denoted as $\mathrm{\rho}$, is crucial for detecting GW events as it quantifies the GW signal’s strength relative to the background noise. An event is deemed detectable if its SNR surpasses a predefined threshold, $\mathrm{\rho_{TH}}$. The optimal SNR for a GW emitted by a binary system is defined by the equation [106, 107, 108]
$\mathrm{\rho^{2}=4\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{|\tilde{h}(f)|^{2}}{S_{n}(f)}\,df},$ (3.3)
where $\mathrm{\tilde{h}(f)}$ represents the Fourier transform of the GWs strain, while $\mathrm{S_{n}(f)}$ denotes the power spectral density of the detector noise.
Figure 3: Noise curve for the LISA showing the square root of the power spectral density, $\mathrm{\sqrt{S_{n}}}$, as a function of frequency. The sensitivity is depicted in units of $\mathrm{1/\sqrt{Hz}}$.
For $\mathrm{S_{n}(f)}$, we adopt the semi-analytical forms from [109], as shown in Figure3, and $\mathrm{f_{min}}$ and $\mathrm{f_{max}}$ are the boundaries for the frequency integral. In our study, for the computation of the SNR, we set $\mathrm{f_{max}}$ to the frequency of the innermost stable circular orbit ($\mathrm{f_{ISCO}}$) and $\mathrm{f_{min}}$ to the frequency one year prior to $\mathrm{f_{ISCO}}$, considering our observation period is one year. We do not include the Galactic background noise in the power spectral density. The Galactic background can have an impact on the overall noise, which in turn affects the number of detectable events.
Specifically, $\mathrm{\tilde{h}(f)}$ is expressed as
$\mathrm{\tilde{h}(f)=h(f)Q},$ (3.4)
with $\mathrm{h(f)}$ denoting the strain amplitude
$\mathrm{h(f)=\sqrt{\frac{5\eta}{24}}\frac{(GM_{\text{chirp}})^{5/6}}{D_{L}\pi^{2/3}c^{3/2}}f^{-7/6}e^{-i\psi(f)}},$ (3.5)
where $\mathrm{M_{chirp}}$ is the chirp mass, $\mathrm{\eta}$ is the mass ratio, $\mathrm{D_{L}}$ is the luminosity distance to the source, $\mathrm{c}$ is the speed of light, and $\mathrm{\psi(f)}$ is the phase of the waveform [110]. The factor $\mathrm{Q}$ depends on the inclination angle and the detection antenna functions $\mathrm{F_{+}}$ and $\mathrm{F_{\times}}$
$\mathrm{Q=\left[F_{+}\left(\frac{1+\cos^{2}(\iota)}{2}\right)+\iota F_{\times}\cos(\iota)\right]}.$ (3.6)
@@
$\mathrm{F_{+}}$ $=\mathrm{\frac{1}{2}(1+\cos^{2}\theta)\cos 2\phi\cos 2\psi-\cos\theta\sin 2\phi\sin 2\psi},$ (3.7)
$\mathrm{F_{\times}}$ $=\mathrm{\frac{1}{2}(1+\cos^{2}\theta)\cos 2\phi\sin 2\psi+\cos\theta\sin 2\phi\cos 2\psi},$ (3.8)
@@
where $\mathrm{\theta}$ and $\mathrm{\phi}$ determine the source’s location in the sky, and $\psi$ relates to the orientation of the binary system with respect to the detector [111]. Consequently, the SNR formula becomes
$\mathrm{\rho=\frac{\Theta}{4}\left[4\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{h(f)^{2}}{S_{n}(f)}df\right]^{1/2}},$ (3.9)
@@
where $\mathrm{\Theta^{2}\equiv 4\left(F_{+}^{2}(1+\cos^{2}i)^{2}+4F_{\times}^{2}\cos^{2}i\right)}$. According to [112], averaging over many binaries, the inclination angle, and sky positions, $\mathrm{\Theta}$ follows the distribution
$\mathrm{P(\Theta)}=\begin{cases}\mathrm{5\Theta(4-\Theta)^{3}/256}&\text{if }\mathrm{0<\Theta<4},\\ \mathrm{0}&\text{otherwise}.\end{cases}$ (3.10)
@@
Figure 4: The figure presents the total number of injected GW events and the number of detectable GW events as a function of redshift for three different models: Light Seed, Heavy Seed, and Heavy Seed+delay. The total observation periods are indicated in parentheses next to the total number of injected events. For the detectable events, the threshold SNR is mentioned in parentheses next to the number of detectable events, assuming that EM counterparts are detected for $\mathrm{75\%}$ of the events.
@@
Figure 5: In this figure, we present a scatter plot illustrating the total mass of the two components of the binary system against the redshift of the selected events, with a threshold SNR of 50. The plot includes data for three different models: Light Seed, Heavy Seed, and Heavy Seed+delay, represented by square, triangle, and circular markers, respectively. The observations span 4 years for the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for $\mathrm{75\%}$ of the events. The color of each point represents the SNR of the event, providing an additional dimension of information.
@@
In this study, we investigate various merger and mass models, recognizing that our understanding of these models is still incomplete. By examining different scenarios, we demonstrate how different assumptions about the formation and evolution of MBBHs impact the measurements of $\mathrm{\mathcal{F}(z)}$ using LISA sources. To estimate the total number of events, we integrate using Equation 3.1 and determine the number of merging events within each redshift bin, selecting a bin size of $\mathrm{\Delta z=0.4}$. Figure 4 displays the number of events binned by redshift for the Light Seed, Heavy Seed, and Heavy Seed + Delay models. The total number of events for each model, along with the observation years, is summarized in Table 2. For each event, we use the inverse transform method to derive the total mass of the binary system and the parameter $\mathrm{\Theta}$ from their probability distributions. We then generate an equal-mass binary system and calculate the SNR using Equation 3.9.
The events with an SNR exceeding the threshold are selected. Since we only consider a mass ratio $\mathrm{q}$ of 1 when generating the binary sources, we retain $\mathrm{75\%}$ of the events after the SNR cutoff. The number of events selected if only $\mathrm{50\%}$ are retained is also noted and listed in Table 2. In Sec. A, we have also shown the results for different fractions of bright standard sirens detectable from LISA. For comparison, in [102], it was found that, for a mission duration of 5 years with an 80% duty cycle up to a redshift of 10 and an SNR cutoff of 10, there were a few hundred events for both light seed and heavy seed scenarios, and about 30 events for the heavy seed with delay scenario. Here, we consider a more pessimistic choice for SNR and redshift when selecting EM bright sources, as we focus only on high SNR golden events for testing GR (see Sec. 7 and appendix A). The detection probability for the Light Seed model is lower compared to the other two models because its mass distribution peaks at much lower values, resulting in reduced signal strength and, consequently, lower SNR. Figure 5 illustrates the total mass of the two components of the binary system as a function of the redshift of the selected events, with an SNR threshold of fifty.
Model $T_{obs}$ (years) Total Events Selected (75%) Selected (50%) 1 Event Detectability
LS 4 158 73 49 6.12%
HS 4 147 98 66 4.54%
HSD 10 49 31 21 14.28%
Table 2: Number of total and detected events for different merger models Light Seed (LS), Heavy Seed (HS), and Heavy Seed+Delay (HSD), with a threshold SNR of 50. The table includes results based on 75% and 50% selection criteria, as well as the percentage of detectability of a single event at each from $z=1$ to $z=6$ (1 event detectability).